# NeRF-Supervised Feature Point Detection and Description

### v1

@@ 4.3 Implementation details
For all experiments, we present results using our own training pipeline for both SuperPoint-PrP and SiLK-PrP. Additionally, we report results for the baseline SuperPoint[6] model using our own reproduced implementation.
To mitigate misprojection errors during the PrP process, we employ a random sampling approach for selecting image $I^{{}^{\prime}}$ from a subset of the training data. This subset is constrained within a range determined by a lower limit, $\lambda_{l}$, and an upper limit, $\lambda_{u}$, based on the input image $I$. Specifically, we set $\lambda_{l}$ to 70 and $\lambda_{u}$ to 150, ensuring a balance between viewpoint changes while avoiding projection errors. Moreover, the depth rectification process threshold $\epsilon$ was set as $3e^{-2}$ meters.
Lastly, for SuperPoint-PrP and SiLK-PrP training, we apply the same photometric augmentation applied on the baseline models [6, 14] such as random brightness and contrast, Gaussian noise, motion blur, but do not apply homographic augmentation such as rotation, scaling or translation. Lastly, all models are trained using PyTorch [38] and a single NVIDIA RTX 3090 Ti GPU.
@@
To produce the MagicPoint-PrP model, we further train the base MagicPoint model through two rounds of Projective Adaptation on the NeRF dataset, each lasting 30,000 iterations. Finally, the SuperPoint-PrP model is obtained by training both the detector and descriptor heads of MagicPoint-PrP once again on the NeRF dataset for a total of 300,000 iterations. The training process is performed using the baseline SuperPoint model hyperparameters with ADAM optimiser, a learning rate of $1e^{-3}$ and betas set to (0.9, 0.999). The batch size is set at 32 for the MagicPoint-PrP model and 2 for the SuperPoint-PrP model.
@@ 5 Experiments
In this section, we assess the performance of the PrP-trained interest point detector and descriptor models in comparison to their corresponding baseline models. The evaluation encompasses accessing the interest point detection quality and homography estimation on the HPatches dataset [39]. Additionally, on ScanNet [18], we evaluate the indoor relative pose estimation and point cloud registration capabilities of each model. Lastly, we evaluate the outdoor relative pose estimation of each model using the YFCC100M [40] dataset. For all evaluations, we perform feature matching using mutual nearest neighbour. Through the aforementioned experiments, we analyse the following:
  • Our NeRF dataset is approximately $\frac{1}{30}$ the size of the MS-COCO dataset [17], and $\frac{1}{250}$ the size of ScanNet’s [18]. Despite the relatively small scale of our training dataset, we analyse the generalisability of the NeRF-trained models on real-world data and numerous conditions across multiple datasets. Surprisingly, the PrP-models trained on our NeRF datasets consistently demonstrate comparable results in most evaluations. This poses the question of the necessity for large open-source datasets in training learning-based detectors and descriptors.
  • Unlike [17, 18, 19, 41], our NeRF dataset fully consists of synthetic images of real scenes. Thus, we assess the interest point detection quality of the NeRF-trained detectors relative to their corresponding baseline models. Conducted experiments revealed that the quality of interest points generated by NeRF-trained models remained consistently high, showing no degradation compared to the baseline models.
  • Lastly, as we explicitly render each scene in our NeRF dataset from various viewpoints, simulating real camera motion instead of relying on homography to mimic camera motions, we explore whether models supervised by the PrP process during training demonstrate improvement in their relative pose estimation capabilities.
@@ 5.1 Homography Estimation
Following [23, 6, 22, 14, 7] we evaluate homography estimation on the HPatches dataset [39]. The HPatches dataset consist of 57 scenes with varying illumination, and 59 scenes with large viewpoint variation. Each scene contains 6 images, related by ground truth homographies.
@@
We follow [23, 14] are resize the shorter image side to a dimension of 480. To evaluate interest point detection quality and performance, we report the repeatability metric as in [6, 14]. Moreover, we evaluate the keypoint descriptors, we report the mean matching accuracy (MMA) i.e. the ratio of matches where reprojection error is below a specified threshold [20]; and matching score i.e. The ratio between correctly matched points and total number of detected points. Lastly, we use OpenCV with RANSAC algorithim to estimate the homography between an image pair, and report the homography estimation accuracy and area under the curve (AUC) error over all image pairs. We keep the top 1k detected interest points for SuperPoint and 10k for SiLK.
@@
We assess SuperPoint-PrP by comparing it to the baseline SuperPoint model [6]. Additionally, we introduce a hybrid variant of SuperPoint-PrP denoted as "SP-PrP-Hyb". This variant incorporates the MagicPoint model, trained through two rounds of Homographic Adaptation, followed by training SuperPoint model on the NeRF dataset using our PrP process and Projective Adaptation. Furthermore, we present two versions of PrP SiLK models; the first, SiLK-PrP, is trained without rotation or scaling augmentation as detailed in Sec. 4.1. The second, SiLK-PrP-Aug, includes rotation and scaling during training.
@@
As seen in Tab. 1, the PrP trained models are outperformed by their baseline counterparts [14, 6] in all of the HPatches metrics. Nonetheless, this outcome is expected as both SiLK-PrP and SuperPoint-PrP are not rotation or scale invariant, in contrast to their corresponding baseline models. On the other hand, despite SiLK-PrP-Aug’s efforts to incorporate in-place augmentation for rotation and scale invariance (see Fig. 4), it falls short compared to the SiLK model.
@@ 5.2 Relative Pose Estimation
For evaluating the relative pose estimation on indoor scenes, we leverage the ScanNet dataset [18]. We employ the 1500 test image pairs provided by [22].
To assess pose estimation on outdoor scenes, we employ the YFCC100M dataset [40]. We utilise the YFCC100M test split introduced by [22]. This split samples 1000 image pairs from four scenes each, thus, the total number of image pairs for the outdoor pose evaluation is 4000 image pairs.
@@
Additionally, for outdoor relative pose evaluation, we resize the YFCC100M images so that the longest dimension is equal to 1200. We use the top 2k predicted points for SuperPoint and 20k for SiLK.
@@
To address this issue, we refined the indoor relative pose estimation assessment. For scenes where $||t_{GT}||$ is below the threshold $\epsilon$ set at 0.15, we exclusively report the pose error AUC based on angular rotation error. Furthermore, for scenes where the threshold $\epsilon$ exceeds 0.15, we report the pose error AUC as the maximum value between angular rotation error and angular translation error. As presented in Tab. 4, it is evident that all models excel in estimating relative rotation in scenes characterised by minimal translation between viewpoints. Notably, the PrP models exhibit a slight performance advantage over their baseline counterparts in this specific scenario. Additionally, in scenes where the ground truth relative translation exceeds the defined threshold, further improvement can be observed for the PrP models compared to results reported in Tab. 3.
[Uncaptioned image] Figure 5: Instability in Angular Translation Error. Density plot illustrating how the angular translation error is unstable in situations where the ground truth relative translation between two camera viewpoints is minimal.
Table 3: Indoor Relative Pose Estimation Error AUC. The models trained using the PrP process outperform their respective baseline models [14, 6] in accurately estimating relative poses within indoor scenes for all reported angular thresholds. Pose Est. AUC @$5^{\circ}$uparrow @$10^{\circ}$uparrow @$20^{\circ}$uparrow SP 4.95 11.90 21.67 SP-PrP-Hyb 5.78 13.52 24.74 SP-PrP 5.20 12.53 22.82 SiLK 7.18 16.14 27.84 SiLK-PrP 9.42 20.17 32.31 SiLK-PrP-Aug 9.60 20.75 34.24
@@
Similar to results obtained in Tab. 3, the PrP models surpass their corresponding baseline models in relative pose estimation for outdoor scenes, as observed by results shown in Tab. 5. The results reported in Tabs. 3 to 5 collectively highlight that models subjected to PrP supervision during training demonstrate significant improvement in their relative pose estimation capability.
Table 5: Outdoor Relative Pose Estimation Error AUC. The PrP trained models exhibit superior performance compared to their baseline counterparts [14, 6] in estimating relative poses for outdoor scenes, similar to observed results in Tab. 3
@@ 5.3 Pairwise Point Cloud Registration
The point cloud registration evaluation is conducted by resizing the ScanNet [18] images to dimensions of [128 $\times$ 128]. Given a pair of RGB-D images, a 6-DOF pose is estimated that best aligns the first image to the second image. The metrics reported during evaluation are angular rotation and translation errors in degrees and centimetres respectively. Moreover, the chamfer distance is also reported between the ground-truth point cloud and the reconstructed point cloud in centimetres. For all three reported metrics the accuracy, mean and median errors over the entire dataset are reported at different thresholds [44].
@@
Unlike previous evaluations (Secs. 5.1 and 5.2) where sparse feature matching was employed, we follow [44, 14] and perform the point cloud registration using dense feature matching. The ratio-test is utilised to find correspondences and alignment is seen as Procustes problem solved by a weighted Kabsch’s algorithim. [45, 46, 14].
@@
It is essential to highlight that SiLK is trained on a relatively small image size of [164x164], whereas the NeRF dataset comprises larger images [640x480]. Since the point cloud registration is conducted on an image size of [128x128], it is expected that SiLK outperforms SiLK-PrP at lower image dimensions, however, results also indicate that the SiLK PrP trained models could detect high quality interest points at low resolutions.
@@ 6 Conclusion
This paper introduces a novel approach to supervise learning-based feature point detectors and descriptors, leveraging perspective projective geometry on synthesised NeRF data. Despite the fact that our proposed dataset is exclusively composed of synthetic images rather than real RGB images, and is comparatively smaller than large open-source datasets, results indicated there was no observed decrease in generalisability or feature point detection quality. As expected our models generally outperform homography-trained baselines on multi-view benchmarks with highly non-planar scenes, while slightly underperforming on a homography estimation benchmark. The larger potential for further advancements lies in improving the quality of training data with advances in neural rendering that can generate higher quality synthesised images devoid of artefacts, and most importantly more precise depth maps to avoid misprojection errors.

### v2

@@ 4.3 Implementation details
For all experiments, we present results using our training pipeline for both SuperPoint-PrP and SiLK-PrP. Additionally, both SuperPoint-PrP and SiLK-PrP were trained exclusively on the NeRF dataset. Furthermore, we report results for the baseline SuperPoint[6] model using our own reproduced implementation, which closely follows the training hyperparameters of the public version of SuperPoint [39].
To mitigate misprojection errors during the PrP process, we employ a random sampling approach for selecting image $I^{{}^{\prime}}$ from a subset of the training data. This subset is constrained within a range determined by a lower limit, $\lambda_{l}$, and an upper limit, $\lambda_{u}$, based on the input image $I$. Specifically, we set $\lambda_{l}$ to 70 and $\lambda_{u}$ to 150, ensuring a balance between viewpoint changes while avoiding projection errors. Moreover, the depth window estimation process threshold $\epsilon_{d}$ was set as $3e^{-2}$ meters.
Additionally, for SuperPoint-PrP and SiLK-PrP training, we apply the same photometric augmentation applied on the baseline models [6, 14] such as random brightness and contrast, Gaussian noise, motion blur, but do not apply homographic augmentation such as rotation, scaling or translation. Lastly, all models are trained using PyTorch [40] and a single NVIDIA RTX 3090 Ti GPU.
@@
To produce the MagicPoint-PrP model, we further train the base MagicPoint model through two rounds of Projective Adaptation on the NeRF dataset, each lasting 30,000 iterations. Finally, the SuperPoint-PrP model is obtained by training both the detector and descriptor heads of MagicPoint-PrP once again on the NeRF dataset for a total of 300,000 iterations. The training process is performed using the baseline SuperPoint model hyperparameters with ADAM optimiser, a learning rate of $1e^{-3}$ and betas set to (0.9, 0.999). The batch size is set at 32 for the MagicPoint-PrP model and 2 for the SuperPoint-PrP model. Lastly, for the Baseline SuperPoint model, the default and optimum homography-induced correspondence threshold $\epsilon_{s}$ in Eq. 7 is 8. However, experiments showed that the optimum PrP-induced correspondence threshold $\epsilon_{s}$ in Eq. 10 is 4.
@@ 5 Experiments
In this section, we assess the performance of PrP-supervised models compared to their corresponding baselines. Evaluations include interest point detection quality and homography estimation on HPatches [41], indoor relative pose estimation and point cloud registration on ScanNet [15], and outdoor relative pose estimation on YFCC100M and MegaDepth [16, 17]. For all evaluations, we use mutual nearest neighbour for feature matching. Through these experiments, we analyse the following:
  • Our NeRF dataset is approximately $\frac{1}{30}$ the size of the MS-COCO dataset [20], and $\frac{1}{250}$ the size of ScanNet’s [15]. Despite the relatively small scale of our training dataset, we analyse the generalisability of the NeRF-trained models on real-world data and various conditions across multiple datasets. Surprisingly, the PrP models trained on our NeRF datasets consistently demonstrate comparable results in most evaluations. This raises the question of the necessity for large open-source datasets in training learning-based detectors and descriptors.
  • Unlike [20, 15, 17, 42], our NeRF dataset consists entirely of synthetic images. We assess the interest point detection quality of the NeRF-trained detectors against their baselines. Experiments revealed that the quality of interest points generated by NeRF-trained models remained consistently high, showing no degradation compared to the baseline models.
  • As we explicitly render each scene in our NeRF dataset from various viewpoints, simulating real camera motion instead of relying on homography to mimic camera motions, we explore whether models supervised by the PrP process demonstrate improvement in their relative pose estimation capabilities. Additionally, we assess whether PrP-supervised models require fewer iterations during training compared to the homography-supervised baselines. Results show that not only do the PrP-supervised models exhibit enhancement in relative pose estimation but also converge faster (i.e. SiLK-PrP converged in approximately 27K iterations compared to 50K for SiLK).
@@ 5.1 Homography Estimation
Following [24, 6, 23, 14, 7] we evaluate homography estimation on the HPatches dataset [41]. The HPatches dataset consists of 57 scenes with varying illumination and 59 scenes with large viewpoint variations. Each scene contains 6 images, related by ground truth homographies.
@@
We follow [24, 14] and resize the shorter image side to a dimension of 480. To evaluate interest point detection quality and performance, we report the repeatability metric as in [6, 14]. Moreover, we evaluate the keypoint descriptors, we report the mean matching accuracy (MMA) i.e. the ratio of matches where reprojection error is below a specified threshold [21], and matching score i.e. The ratio between correctly matched points and a total number of detected points. Lastly, we use OpenCV’s RANSAC algorithm to estimate the homography between an image pair and report the homography estimation accuracy and area under the curve (AUC) error over all image pairs. We keep the top 1k detected interest points for SuperPoint and 10k for SiLK.
@@
We assess SuperPoint-PrP by comparing it to the baseline SuperPoint model [6]. Additionally, we introduce a hybrid variant of SuperPoint-PrP denoted as "SP-PrP-Hyb". This variant incorporates the MagicPoint model, trained through two rounds of Homographic Adaptation, followed by training the SuperPoint model on the NeRF dataset using our PrP process and Projective Adaptation. Furthermore, we present two versions of PrP SiLK models; the first, SiLK-PrP, is trained without rotation or scaling augmentation as detailed in Sec. 4.1. The second, SiLK-PrP-Aug, includes rotation and scaling augmentations during training.
@@
As seen in Tab. 1, the PrP-supervised models are outperformed by their baseline counterparts [14, 6] in all of the HPatches metrics. Nonetheless, this outcome is expected as both SiLK-PrP and SuperPoint-PrP are not rotation or scale-invariant, in contrast to their corresponding baseline models. On the other hand, despite SiLK-PrP-Aug’s efforts to incorporate in-place augmentation for rotation and scale invariance (see Fig. 3), it falls short compared to the SiLK model.
@@ 5.2 Relative Pose Estimation
We leverage the ScanNet dataset [15] to evaluate the relative pose estimation on indoor scenes. We employ the 1500 test image pairs provided by [23].
We employ the YFCC100M [16] and MegaDepth [17] datasets to assess pose estimation on outdoor scenes. We utilise the YFCC100M test split introduced by [23], this split samples 1000 image pairs from four scenes each, thus, the total number of image pairs for the outdoor pose evaluation is 4000 image pairs. Additionally, we use the MegaDepth1500 test split introduced by [21, 24]. This test split comprises 1500 image pairs from two phototourism scenes: “Sacre Coeur” and “St. Peter’s Square”.
@@
Additionally, for outdoor relative pose evaluation, we resize the YFCC100M and MegaDepth1500 images so that the longest dimension is equal to 1200 and 1600 respectively. We use the top 2k predicted points for SuperPoint and 20k for SiLK.
@@
To address this issue, we refined the indoor relative pose estimation assessment. For scenes where $||t_{GT}||$ is below the threshold $\epsilon$ set at 0.15, we exclusively report the pose error AUC based on angular rotation error. Furthermore, for scenes where the threshold $\epsilon$ exceeds 0.15, we report the pose error AUC as the maximum value between angular rotation and translation errors. As presented in Tab. 4, it is evident that all models excel in estimating relative rotation in scenes characterised by minimal translation between viewpoints. Notably, the PrP models exhibit a slight performance advantage over their baseline counterparts in this specific scenario. Additionally, in scenes where the ground truth relative translation exceeds the defined threshold, further improvement can be observed for the PrP models compared to results reported in Tab. 3.
Figure 4: Instability in Angular Translation Error. Density plot illustrating how the angular translation error is unstable in situations where the ground truth relative translation between two camera viewpoints is minimal.
Table 3: Indoor Relative Pose Estimation Error AUC. The models trained using the PrP process outperform their respective baseline models [14, 6] in accurately estimating relative poses within indoor scenes for all reported angular thresholds.
Pose Est. AUC @$5^{\circ}$uparrow @$10^{\circ}$uparrow @$20^{\circ}$uparrow
SP 7.56 17.76 30.43
SP-PrP-Hyb 7.82 17.85 31.08
SP-PrP 7.73 17.72 30.71
SiLK 7.18 16.14 27.84
SiLK-PrP 9.42 20.17 32.31
SiLK-PrP-Aug 9.60 20.75 34.24
@@
Similar to results obtained in Tab. 3, the PrP models surpass their corresponding baseline models in relative pose estimation for outdoor scenes, as observed by results shown in Tabs. 5 and 6. The results reported in Tabs. 3 to 6 collectively highlight that models subjected to PrP supervision during training demonstrate significant improvement in their relative pose estimation capability.
Table 5: YFCC Relative Pose Estimation
Error AUC.
@@
Table 6: MegaDepth Relative Pose Estimation
Error AUC.
Pose Est. AUC @$5^{\circ}$uparrow @$10^{\circ}$uparrow @$20^{\circ}$uparrow
SP 25.21 39.42 52.11
SP-PrP-Hyb 27.34 42.64 56.03
SP-PrP 26.17 41.53 55.87
SiLK 22.70 35.32 48.08
SiLK-PrP 25.91 35.50 46.41
SiLK-PrP-Aug 24.79 34.38 45.00
@@ 5.3 Pairwise Point Cloud Registration
The point cloud registration evaluation is conducted by resizing the ScanNet [15] images to dimensions of [128 $\times$ 128]. Given a pair of RGB-D images, a 6-DOF pose is estimated that best aligns the first image to the second image. The metrics reported during the evaluation are angular rotation and translation errors in degrees and centimetres respectively. Moreover, the chamfer distance is also reported between the ground-truth point cloud and the reconstructed point cloud in centimetres. For all three reported metrics the accuracy, mean and median errors over the entire dataset are reported at different thresholds [45].
@@
Unlike previous evaluations (Secs. 5.1 and 5.2) where sparse feature matching was employed, we follow [45, 14] and perform the point cloud registration using dense feature matching. The ratio test is utilised to find correspondences and alignment is seen as a Procustes problem solved by a weighted Kabsch’s algorithm. [46, 47, 14].
@@
It is essential to highlight that SiLK is trained on smaller images [164x164], while the NeRF dataset is comprised of larger images [640x480]. Since the point cloud registration is conducted on an image size of [128x128], SiLK may outperform SiLK-PrP at lower resolutions, however, the SiLK-PrP models still detect high-quality interest points at low resolutions.
@@ 6 Conclusion
This paper presented a novel approach to supervise feature point detectors and descriptors using perspective projective geometry on synthetic data. Despite using a smaller dataset, our PrP approach achieves similar or better performance without compromising generalisability. Moreover, we noticed that the PrP-supervised models converge faster and outperform homography-supervised models in non-planar multi-view benchmarks, though, they slightly lag in homography estimation. We emphasise the main advantage of the PrP approach is not only benchmark improvements, but also its ability to train on any custom multi-view synthetic dataset with less data and less training time. The larger potential for further advancements depends on enhancing neural rendering to produce higher-quality synthetic images and more accurate depth maps, reducing artefacts and misprojection errors.